OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomores tickets totaled $410. How many sophomores bought tickets? A. 40 B. 90 C. 110 D. 120 I believe the answer is C, is that right?

OpenStudy (lgbasallote):

let x = freshmen y = sophomores x+y = 150 2x + 3y = 410 -2x -2y = -300 2x + 3y = 410 y = 110 x = 40

OpenStudy (lgbasallote):

ys 110 sophomores...you're right

OpenStudy (anonymous):

Thank you! :)